OpenStudy (suenahelpz):

A pair of equations is shown below. x + y = 2 y = 1/2x + 5 If the two equations are graphed, at what point do the lines representing the two equations intersect? (4, −2) (−2, 4) (2, 5) (5, −2) My guess would be A..

OpenStudy (mathmale):

Suggestion: Use parentheses so that there's no doubt regarding what you mean by y = 1/2x + 5 . If your 1/2 is the slope, then it's essential that you enclose that 1/2 in parentheses. Otherwise your y = 1/2x + 5 could be interpreted as y = 1/(2x) + 5, which is an entirely different function.

OpenStudy (mathmale):

Have you tried graphing your version of this function? If you do this accurately, it should be straightforward to determine the point of intersection.

OpenStudy (triciaal):

at the intersection (x,y) is the same for both equations write the expression for x or y from equation 1 and set it equal to the same expression in equation 2 solve for the one unknown in that equation use that value in the expression to calculate the value for the other variable

OpenStudy (suenahelpz):

I have tried graphing it @mathmale, but it ends terribly wrong since I am very confused as to what to graph.

OpenStudy (triciaal):